### v2

@@ 4 Manifolds of absolutely continuous functions
and the homeomorphism
@@ 6 Semiregularity of $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$
Let $n,m,r\in\mathbb{N}$. We set the Fréchet space $C(\mathbb{R}^{n},\mathbb{R}^{m})$ of continuous functions $\phi:\mathbb{R}^{n}\to\mathbb{R}^{m}$, whose topology is generated by the famility of seminorms
@@
$C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{m}):=\{\phi\in C^{r}(\mathbb{R}^{n},\mathbb{R}^{m}):\phi|_{\mathbb{R}^{n}\setminus K}=0\}$
@@
$D:C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}))\to C_{K}^{r-1}(\mathbb{R}^{n},\mathbb{R}^{n\times n}),\quad f\mapsto Df:=f^{\prime}$
@@
$\text{Diff}_{K}^{r}(\mathbb{R}^{n}):=\{\phi\in\text{Diff}^{r}(\mathbb{R}^{n}):\phi-\text{id}_{\mathbb{R}^{n}}\in C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\}.$
@@
$\Phi:\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})\to\mathcal{V}_{K},\quad\phi\mapsto\phi-\text{id}_{\mathbb{R}^{n}}$
is a global chart for $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$, turning it into a right half-Lie group modeled on the Banach space $C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})$ (See [12, Proposition 14.6]). On the set $\mathcal{V}_{K}$ we define a group multiplication by
@@
($\left(\{0\}\times C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\right)\times\mathcal{V}_{K}\to T\mathcal{V}_{K},\quad((0,\varphi),\psi)\mapsto d\rho_{\psi}(0,\varphi)=\varphi\circ(id_{\mathbb{R}^{n}}+\psi)$
@@
Since the evaluation map $\varepsilon:C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ is $C^{r,\infty}$, by Lemma 5.12 the map
@@
for $t\in[0,1]$. Then $T$ is $C^{\infty,\infty,r}$.
@@
Hence $T([\gamma],x,\cdot)$ is a $C^{r}$-map and a $L$-contraction. By Banach’s Fixed Point Theorem, there exists $\zeta_{[\gamma],x}\in C([0,1],\mathbb{R}^{n})$ such that
@@
$S:B\to C([0,1],C^{r}(\mathbb{R}^{n},\mathbb{R}^{n})),\quad[\gamma]\mapsto F_{4}([\gamma])-\text{I}.$
@@ 7 $L^{p}$-Semiregularity of $\text{Diff}^{\text{ }r}(M)$
turning $\text{Diff}^{\text{ }r}(M)$ in a Banach manifold. Moreover, under the composition the manifold $\text{Diff}^{\text{ }r}(M)$ becomes a right half-Lie group (see e.g. [12]). We see that for the neutral element $e=\text{id}_{M}$, the tangent space $T_{e}\text{Diff}^{\text{ }r}(M)$ coincide with space of $C^{r}$-vector fields $X:M\to TM$, denoted by $\mathcal{X}^{r}(M)$.
@@
$\mathcal{X}^{r}(M)\times\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})\to T\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n}),\quad(X,\phi)\mapsto X.\phi=X\circ\phi$
@@
$\dot{\eta}=\gamma.\eta,\quad\eta(0)=\text{id}_{M}.$
@@
Let $(E,\lVert\cdot\lVert_{E})$ be a normed space and $U\subseteq E$ be a subset. We say that a function $f:[a,b]\times U\to E$ satisfies a $L^{1}$-Lipschitz condition if there exists a measurable function $g\in L^{1}([a,b],\mathbb{R})$ such that
$\text{Lip}(f(t,\cdot))\leq g(t),\quad t\in[a,b]$
Where $\text{Lip}(f(t,\cdot))\in[0,\infty]$ denote the infimum of all Lipschitz constant for $f(t,\cdot):U\to E$.
@@
satisfying $\tau(t_{0})=\eta(t_{0})$ for some $t_{0}\in[a,b]$, then $\tau=\eta$ [10, Proposition 10.5]. If the initial value problem
@@
The following result follows the same steps as the proof of the $L^{1}$-regularity of $\text{Diff}_{c}^{\infty}(M)$ proved in [10, Section 11]. We recommend to the reader to read the reference to see some steps in detail.
@@
$X_{i}:=d\psi_{i}\circ X\circ\kappa_{i}^{-1}:B_{5}(0)\to\mathbb{R}^{n}.$
@@
$\rho_{\ell}:\Gamma_{C^{r}}(\phi)\to\prod_{i=1}^{n}C^{r}(B_{\ell}(0),\mathbb{R}^{n}),\quad X\mapsto\rho_{\ell}(X):=(X_{i}|_{B_{\ell}(0)})_{i=1}^{n}$
@@
$R_{5}:=L^{p}([0,1],\rho_{5}):L^{p}([0,1],\mathcal{X}^{r}(M))\to\prod_{i=1}^{n}L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{n})),\quad\gamma\mapsto\rho_{5}\circ\gamma$
@@
$R_{1}:=C([0,1],\rho_{1}):C([0,1],\mathcal{X}^{r}(M))\to\prod_{i=1}^{n}C([0,1],C^{r}(B_{1}(0),\mathbb{R}^{n})),\quad\gamma\mapsto\rho_{1}\circ\gamma.$
Before study the equation for $[\gamma]\in L^{p}([0,1],\mathcal{X}^{r}(M))$, the embedding $R_{1}$ allow us to study it first for
$[\gamma]\in L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{n})).$
We consider the continuous seminorm $q:C^{r}(B_{5}(0),\mathbb{R}^{n})\to[0,\infty)$ by
@@
$B_{L}:=\{\gamma\in L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{n})):\lVert\gamma\lVert_{L^{p},q}<L\}.$
Just like the case for $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$, we need a $AC_{L^{p}}$-Carathéodory solution $\zeta:[0,1]\to B_{4}(0)$ that verifies the integral equation
@@
is $C^{r}$ (see [10, Lemma 6.2]). By the exponential law, we have
@@
$\lfloor\overline{B_{2}(0)},B_{4}(0)\rfloor_{r}:=\{\varphi\in C^{r}(B_{3}(0),\mathbb{R}^{n}):\varphi(\overline{B_{2}(0)})\subseteq B_{4}(0)\}.$
@@
$S:C^{r}(B_{5}(0),\mathbb{R}^{n})\times\lfloor\overline{B_{2}(0)},B_{4}(0)\rfloor_{r}\to C^{r}(B_{2}(0),\mathbb{R}^{n}),\quad(\psi,\varphi)\mapsto\psi\circ\varphi$
@@
$\tilde{S}:L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{n}))\times C([0,1],\lfloor\overline{B_{2}(0)},B_{4}(0)\rfloor_{r})\to L^{p}([0,1],C^{r}(B_{2}(0),\mathbb{R}^{n})),\quad(\alpha,\beta)\mapsto S\circ(\alpha,\beta)$
@@
$S\circ(\gamma,H([\gamma]))=\gamma\left(H(\gamma)\right)=\gamma.H([\gamma])\in L^{p}([0,1],C^{r}(B_{2}(0),\mathbb{R}^{n})).$
This allow us to define the $L^{p}$-absolute continuous function $\tau:[0,1]\to C^{r}(B_{2}(0),\mathbb{R}^{n})$ given by
@@
If we consider the point evaluation map $\varepsilon_{x}:C^{r}(B_{2}(0),\mathbb{R}^{n})\to\mathbb{R}^{n}$, $\varphi\mapsto\varphi(x)$, then for each $x\in B_{2}(0)$ we have
@@
Since the family of maps $(\varepsilon_{x})_{x\in B_{2}(0)}$ separate points in $C^{r}(B_{2}(0),\mathbb{R}^{n})$, we have $\tau=H([\gamma])$.
@@
Let $[\gamma]\in\mathcal{B}_{L}$, with $[\gamma_{i}]\in B_{L}$ for each $i\in\{1,...,n\}$. If $p\in M$, there is a chart $\kappa_{j}:U_{j}\to B_{5}(0)$ such that $p\in\kappa_{j}^{-1}(B_{3}(0))$ for some $j\in\{1,...,n\}$. Let $\zeta_{[\gamma],\kappa_{j}(p)}\in AC_{L^{P}}([0,1],C^{r}(B_{2}(0),B_{4}(0))$ be the solution of the integral equation
@@
Then $\zeta_{[\gamma],\kappa_{j}(p)}$ is a $AC_{L^{p}}$-Carathéodory solution to
@@
$\eta_{[\gamma],p}:[0,1]\to M,\quad\eta_{[\gamma],p}(t):=\kappa_{j}^{-1}\circ\zeta_{[\gamma],\kappa_{j}(p)}(t)$
@@
$y(0)$ $=p.$
Since $\tilde{\gamma}$ satisfies a local $L^{1}$-Lipschitz condition, by Remark 7.3 the solution of this equation is unique and $\eta_{[\gamma],p}$ is well defined. Moreover, $\tilde{\gamma}$ admits a global flow for initial time $t_{0}=0$, given by
$F_{t,0}^{\overline{\gamma}}(p)=\eta_{[\gamma],p}(t)$
for all $t\in[0,1]$, $p\in M$.
@@
Moreover, for an open subset $\mathcal{O}_{i}\subseteq\mathcal{D}_{i}$ we consider the map
$\psi_{i}:=(\text{pr}_{1},\text{exp}_{i})|_{\mathcal{O}_{i}}:\mathcal{O}_{i}\subseteq TB_{5}(0)\to\psi_{i}(\mathcal{O}_{i})\subseteq B_{5}(0)\times B_{4}(0).$
Assuming that
@@
there is exists $s_{i}\in]0,1]$ such that
@@
is a smooth diffeomorphism and there exists a $s_{i}\in]0,1[$, such that $\psi_{i}^{-1}$ restrict to a smooth diffeomorphism of the form
@@
$\mathcal{O}_{i,x}:=\{y\in\mathbb{R}^{n}:(x,y)\in\mathcal{O}_{i}\}.$
@@
is open in $C^{r}(B_{2}(0),\mathbb{R}^{n})$ and the map
$(\theta_{i})_{*}:Z_{i,x}\subseteq C^{r}(B_{2}(0),\mathbb{R}^{n})\to C^{r}(B_{1}(0),\mathbb{R}^{n}),\quad\varphi\mapsto(\theta_{i})_{*}(\varphi):=\theta_{i}\circ(\text{id}_{B_{1}(0)},\varphi)$
is smooth since $\theta$ is smooth [12, Proposition 4.23].
@@
for suitable open 00-neighborhoods $\mathcal{V}_{i}\subseteq C^{r}(B_{1}(0),\mathbb{R}^{n})$.
@@
$\mathcal{P}:=(R_{\text{id}_{M},t})^{-1}\left(\prod_{i=1}^{n}\mathcal{P}_{i}\right)$
@@
$(\theta_{i})_{*}\circ H:B_{L}\subseteq L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{n})\to C([0,1],C^{r}(B_{1}(0),\mathbb{R}^{n})$
@@
Let $p\in M$ with $x:=\kappa_{i}(p)$, then we have
$\exp\circ\theta_{[\gamma]}(t)(p)$ $=\left(\kappa_{i}^{-1}\circ\exp_{i}\circ T\kappa_{i}\right)\circ\theta_{[\gamma]}(t)(\kappa_{i}^{-1}(x))$
@@
$[0,1]\to\prod_{i=1}^{n}C^{r}(B_{1}(0),\mathbb{R}^{n}),\quad t\mapsto\rho_{1}\circ\theta_{[\gamma]}(t)$
But this maps coincide with the map
$[0,1]\to\prod_{i=1}^{n}C^{r}(B_{1}(0),\mathbb{R}^{n}),\quad t\mapsto\left((\theta_{i})_{*}\left(H([\gamma_{i}])\right)\right)_{i=1}^{n}(t).$
is absolutely continuous. Since $(\theta_{i})_{*}$ is smooth and $H([\gamma_{i}])\in AC_{L^{p}}([0,1],C^{r}(B_{2}(0),B_{4}(0)))$, by Lemma 2.16, the composition and hence the product are absolutely continuous. Then we have that
@@
In consequence, Lemma 5.9 and Remark 7.3, the right half-Lie group $\text{Diff}^{\text{ }r}(M)$ is $L^{p}$-semiregular. Since the map

### v4

@@ 3 Manifolds of absolutely continuous functions
which is a homeomorphism onto its image the homeomorphism onto its image (see [37, Proposition 4.13]) and the homeomorphism
@@ 5 Semiregularity of $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$
Let $n,m,r\in\mathbb{N}$. We consider the Fréchet space $C(\mathbb{R}^{n},\mathbb{R}^{m})$ of continuous functions $\phi:\mathbb{R}^{n}\to\mathbb{R}^{m}$, whose topology is generated by the famility of seminorms
@@
$C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{m}):=\{\phi\in C^{r}(\mathbb{R}^{n},\mathbb{R}^{m}):\phi|_{\mathbb{R}^{n}\setminus K}=0\}$ (5.1)
@@
$D:C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\to C_{K}^{r-1}(\mathbb{R}^{n},\mathbb{R}^{n\times n}),\quad f\mapsto Df:=f^{\prime}$ (5.2)
@@
$\text{Diff}_{K}^{r}(\mathbb{R}^{n}):=\{\phi\in\text{Diff}^{r}(\mathbb{R}^{n}):\phi-\text{id}_{\mathbb{R}^{n}}\in C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\}.$ (5.3)
@@
$\Phi:\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})\to\mathcal{V}_{K},\quad\phi\mapsto\phi-\text{id}_{\mathbb{R}^{n}}$ (5.4)
is a global chart for $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$, turning it into a right half-Lie group modeled on the Banach space $C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})$ (See [16, Proposition 14.6]). On the set $\mathcal{V}_{K}$ we define a group multiplication by
@@
($\left(\{0\}\times C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\right)\times\mathcal{V}_{K}\to T\mathcal{V}_{K},\quad((0,\varphi),\psi)\mapsto d\rho_{\psi}(0,\varphi)=\varphi\circ(id_{\mathbb{R}^{n}}+\psi)$ (5.5)
@@
Since the evaluation map $\varepsilon:C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ is $C^{\infty,r}$ in the sense of [1] and hence $C^{r}$. By Lemma 4.12 the map
@@
for $t\in[0,1]$. Then $T$ is $C^{\infty,\infty,r}$ in the sense of [2].
@@
Hence $T([\gamma],x,\cdot)$ is a $C^{r}$-map and an $L$-contraction. By Banach’s Fixed Point Theorem, there exists $\zeta_{[\gamma],x}\in C([0,1],\mathbb{R}^{n})$ such that
@@
$S:B_{L}\to C([0,1],C^{r}(\mathbb{R}^{n},\mathbb{R}^{n})),\quad[\gamma]\mapsto F_{4}([\gamma])-\text{I}.$
@@ 6 $L^{p}$-Semiregularity of $\text{Diff}^{\text{ }r}(M)$
turning $\text{Diff}^{\text{ }r}(M)$ into a Banach manifold. Moreover, under the composition the manifold $\text{Diff}^{\text{ }r}(M)$ becomes a right half-Lie group (see e.g. [16]). We see that for the neutral element $e=\text{id}_{M}$, the tangent space $T_{e}\text{Diff}^{\text{ }r}(M)$ coincide with space of $C^{r}$-vector fields $X:M\to TM$, denoted by $\mathcal{X}^{r}(M)$.
@@
$\mathcal{X}^{r}(M)\times\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})\to T\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n}),\quad(X,\phi)\mapsto X.\phi=X\circ\phi$ (6.1)
@@
$\dot{\eta}(t)$ $=\gamma(t).\eta(t),\quad t\in[0,1]$ (6.2)
$\eta(0)$ $=\text{id}_{M}.$ (6.3)
@@
Let $(E,\lVert\cdot\lVert_{E})$ be a normed space and $U\subseteq E$ be a subset. We say that a function $f:[a,b]\times U\to E$ satisfies an $L^{1}$-Lipschitz condition if there exists a measurable function $g\in L^{1}([a,b],\mathbb{R})$ such that
$\text{Lip}(f(t,\cdot))\leq g(t),\quad t\in[a,b],$
where $\text{Lip}(f(t,\cdot))\in[0,\infty]$ denote the infimum of all Lipschitz constants for $f(t,\cdot):U\to E$.
@@
satisfying $\tau(t_{0})=\eta(t_{0})$ for some $t_{0}\in[a,b]$, then $\tau=\eta$ [14, Proposition 10.5]. If the initial value problem
@@
The following result follows the same steps as the proof of the $L^{1}$-regularity of $\text{Diff}_{c}^{\infty}(M)$ proved in [14, Section 11]. We recommend to the reader to read the reference to see some steps in detail.
@@
$X_{i}:=d\psi_{i}\circ X\circ\kappa_{i}^{-1}:B_{5}(0)\to\mathbb{R}^{m}.$
@@
$\rho_{\ell}:\Gamma_{C^{r}}(\phi)\to\prod_{i=1}^{n}C^{r}(B_{\ell}(0),\mathbb{R}^{m}),\quad X\mapsto\rho_{\ell}(X):=(X_{i}|_{B_{\ell}(0)})_{i=1}^{n}$
@@
$R_{5}:=L^{p}([0,1],\rho_{5}):L^{p}([0,1],\mathcal{X}^{r}(M))\to\prod_{i=1}^{n}L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{m})),\quad\gamma\mapsto\rho_{5}\circ\gamma$
@@
$R_{1}:=C([0,1],\rho_{1}):C([0,1],\mathcal{X}^{r}(M))\to\prod_{i=1}^{n}C([0,1],C^{r}(B_{1}(0),\mathbb{R}^{m})),\quad\gamma\mapsto\rho_{1}\circ\gamma.$
Before we study the equation for $[\gamma]\in L^{p}([0,1],\mathcal{X}^{r}(M))$, the topological embedding $R_{1}$ allows us to study it first for
$[\gamma]\in L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{m})).$
We consider the continuous seminorm $q:C^{r}(B_{5}(0),\mathbb{R}^{m})\to[0,\infty)$ given by
@@
$B_{L}:=\{\gamma\in L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{m})):\lVert\gamma\lVert_{L^{p},q}<L\}.$
Just like the case for $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$, we need an $AC_{L^{p}}$-Carathéodory solution $\zeta:[0,1]\to B_{4}(0)$ that verifies the integral equation
@@
is $C^{r}$ (see [14, Lemma 6.2]). By the exponential law, we have
@@
$\lfloor\overline{B_{2}(0)},B_{4}(0)\rfloor_{r}:=\{\varphi\in C^{r}(B_{3}(0),\mathbb{R}^{m}):\varphi(\overline{B_{2}(0)})\subseteq B_{4}(0)\}.$
@@
$S:C^{r}(B_{5}(0),\mathbb{R}^{n})\times\lfloor\overline{B_{2}(0)},B_{4}(0)\rfloor_{r}\to C^{r}(B_{2}(0),\mathbb{R}^{m}),\quad(\psi,\varphi)\mapsto\psi\circ\varphi$
@@
$\tilde{S}:L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{n}))\times C([0,1],\lfloor\overline{B_{2}(0)},B_{4}(0)\rfloor_{r})$ $\to L^{p}([0,1],C^{r}(B_{2}(0),\mathbb{R}^{m}))$
$(\alpha,\beta)$ $\mapsto S\circ(\alpha,\beta)$
@@
$S\circ(\gamma,H([\gamma]))=\gamma\left(H(\gamma)\right)=\gamma.H([\gamma])\in L^{p}([0,1],C^{r}(B_{2}(0),\mathbb{R}^{m})).$
This allow us to define the $L^{p}$-absolutely continuous function $\tau:[0,1]\to C^{r}(B_{2}(0),\mathbb{R}^{m})$ given by
@@
If we consider the point evaluation map $\varepsilon_{x}:C^{r}(B_{2}(0),\mathbb{R}^{m})\to\mathbb{R}^{m}$, $\varphi\mapsto\varphi(x)$, then for each $x\in B_{2}(0)$ we have
@@
Since the family of maps $(\varepsilon_{x})_{x\in B_{2}(0)}$ separate points in $C^{r}(B_{2}(0),\mathbb{R}^{m})$, we have $\tau=H([\gamma])$.
@@
Let $[\gamma]\in\mathcal{B}_{L}$, with $[\gamma_{i}]\in B_{L}$ for each $i\in\{1,...,n\}$. If $m\in M$, there is a chart $\kappa_{j}:U_{j}\to B_{5}(0)$ such that $m\in\kappa_{j}^{-1}(B_{3}(0))$ for some $j\in\{1,...,n\}$. Let $\zeta_{[\gamma],\kappa_{j}(m)}\in AC_{L^{P}}([0,1],C^{r}(B_{2}(0),B_{4}(0))$ be the solution of the integral equation
@@
Then $\zeta_{[\gamma],\kappa_{j}(m)}$ is a $AC_{L^{p}}$-Carathéodory solution to
@@
$\eta_{[\gamma],m}:[0,1]\to M,\quad\eta_{[\gamma],m}(t):=\kappa_{j}^{-1}\circ\zeta_{[\gamma],\kappa_{j}(m)}(t)$
@@
$y(0)$ $=m.$
Since $\tilde{\gamma}$ satisfies a local $L^{1}$-Lipschitz condition, by Remark 6.3 the solution of this equation is unique and $\eta_{[\gamma],m}$ is well defined. Moreover, $\tilde{\gamma}$ admits a global flow for initial time $t_{0}=0$, given by
$F_{t,0}^{\overline{\gamma}}(m)=\eta_{[\gamma],m}(t)$
for all $t\in[0,1]$, $m\in M$.
@@
There is an open set $\mathcal{O}_{i}\subseteq\mathcal{D}_{i}$ containing $\overline{B_{4}(0)}\times\{0\}$ such that $(\text{pr}_{1},\exp_{i})(Q_{i})$ is open in $B_{5}(0)\times B_{5}(0)$ and the map
$\psi_{i}:=(\text{pr}_{1},\text{exp}_{i})|_{\mathcal{O}_{i}}:\mathcal{O}_{i}\subseteq TB_{5}(0)\to\psi_{i}(\mathcal{O}_{i})\subseteq B_{5}(0)\times B_{4}(0)$
is a $C^{\infty}$-diffeomorphism onto its image. Assuming that
@@
there exists $s_{i}\in]0,1]$ such that
@@
is a smooth diffeomorphism and there exists an $s_{i}\in]0,1[$, such that $\psi_{i}^{-1}$ restrict to a smooth diffeomorphism of the form
@@
$\mathcal{O}_{i,x}:=\{y\in\mathbb{R}^{m}:(x,y)\in\mathcal{O}_{i}\}.$
@@
is open in $C^{r}(B_{2}(0),\mathbb{R}^{m})$ and the map
$(\theta_{i})_{*}:Z_{i,x}\subseteq C^{r}(B_{2}(0),\mathbb{R}^{m})\to C^{r}(B_{1}(0),\mathbb{R}^{m}),\quad\varphi\mapsto(\theta_{i})_{*}(\varphi):=\theta_{i}\circ(\text{id}_{B_{1}(0)},\varphi)$
is smooth since $\theta$ is smooth [16, Proposition 4.23].
@@
for suitable open 00-neighborhoods $\mathcal{V}_{i}\subseteq C^{r}(B_{1}(0),\mathbb{R}^{m})$.
@@
$\mathcal{P}:=R_{5}^{-1}\left(\prod_{i=1}^{n}\mathcal{P}_{i}\right)$
@@
$(\theta_{i})_{*}\circ H:B_{L}\subseteq L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{m})\to C([0,1],C^{r}(B_{1}(0),\mathbb{R}^{m})$
@@
Let $m\in M$ with $x:=\kappa_{i}(m)$, then we have
$\exp\circ\theta_{[\gamma]}(t)(m)$ $=\left(\kappa_{i}^{-1}\circ\exp_{i}\circ T\kappa_{i}\right)\circ\theta_{[\gamma]}(t)(\kappa_{i}^{-1}(x))$
@@
$[0,1]\to\prod_{i=1}^{n}C^{r}(B_{1}(0),\mathbb{R}^{m}),\quad t\mapsto\rho_{1}\circ\theta_{[\gamma]}(t).$
But this map coincides with the map
$[0,1]\to\prod_{i=1}^{n}C^{r}(B_{1}(0),\mathbb{R}^{m}),\quad t\mapsto\left((\theta_{i})_{*}\left(H([\gamma_{i}])\right)\right)_{i=1}^{n}(t)$
which is absolutely continuous. Since $(\theta_{i})_{*}$ is smooth and $H([\gamma_{i}])\in AC_{L^{p}}([0,1],C^{r}(B_{2}(0),B_{4}(0)))$, by Proposition 3.8, the composition and hence the product are absolutely continuous. Then we have that
@@
In consequence, by Lemma 4.9 and Remark 6.3, the right half-Lie group $\text{Diff}^{\text{ }r}(M)$ is $L^{p}$-semiregular. Since the map